A piece of equipment costs a certain factory Rs 6 lakhs. It depreciates in value 15% the first year, 13.5% the next year, 12% the third year and so on. What will be its value at the end of 10 years all percentages applying to the original cost?

Grade:10

3 Answers

Sher Mohammad IIT Delhi
10 years ago
15, 13.5,12 ....,1.5, summation of all this gives 82.5,
6-6*85%=6(1-.85)=6*.15=.9 lakh is the value after 10 years.

sher mohammad
b.tech, iit delhi

Riddhi Roy
10 years ago
but the answer given in RD Sharma is 105000
Suresh Kumar Pillai
6 years ago
15, 13.5, 12 are in A PSo 10th term is 15-9*1.5=1.5Sum of depreciation= 15+13.5+12+...+1.5=82.5Value=17.5%Value =17.5*600000÷100=105000